OpenStudy (anonymous):

What is (cos Θ − cos Θ)^2 + (cos Θ + cos Θ)^2? sin^2 Θ 4cos^2 Θ 8 cos^2 Θ

OpenStudy (astrophysics):

cos theta - cos theta = ?

OpenStudy (astrophysics):

it's like saying x - x, what does that give you?

OpenStudy (anonymous):

I know that part equals out to 0, but then I'm left with (cos theta + cos theta)^2

OpenStudy (anonymous):

I can't figure that out :(

OpenStudy (astrophysics):

Oh but in your case, there's a right answer yay!

OpenStudy (anonymous):

It becomes \[\cos ^{2}\theta + \cos ^{2}\theta\]

zepdrix (zepdrix):

Expand? 0_o But you said x-x = 0 why not apply that thinking to the other one? x+x = 2x

OpenStudy (anonymous):

That's why I'm confused, it can't just equal 2x

zepdrix (zepdrix):

...?

zepdrix (zepdrix):

The other one is x+x

OpenStudy (astrophysics):

the (cos theta+ cos theta)^2 term

OpenStudy (astrophysics):

cos theta + cos theta = 2cos theta, now square it

OpenStudy (anonymous):

Ohhh

OpenStudy (anonymous):

Well then it would be \[4\cos^2 \theta\]

OpenStudy (anonymous):

Well I'm an idiot

OpenStudy (anonymous):

It was simple... -_-

OpenStudy (astrophysics):

Lol, don't worry for some reason I was thinking costheta+ sintheta

OpenStudy (astrophysics):

But in any case if you took long way or just added right away, you'd get same answer

OpenStudy (anonymous):

It's okay, thank you for helping though!

OpenStudy (astrophysics):

Yw :)